2.2 Find the solution of the following couple of differential equations: x^ prime =4x-y and y^ prime =-4x+4y When t = 0, x(0) = 1 and y(0) = 1 .


1
Expert's answer
2021-02-24T07:36:07-0500

"\begin{cases}
\frac{dx}{dt}=4x-y\\
\frac{dy}{dt}=-4x+4y
\end{cases}"


Characteristic equation:

"\begin{vmatrix}
4-\lambda & -1\\
-4 & 4-\lambda
\end{vmatrix}=(4-\lambda)^2-4="

"=(4-\lambda-2)(4-\lambda+2)=(2-\lambda)(6-\lambda)"

Eigen values:

"\lambda_1=2"

"\lambda_2=6"

Then we find eigen vectors:


for "\lambda_1" :

"\begin{pmatrix}
2 & -1\\
-4 & 2
\end{pmatrix}
\begin{pmatrix}
x_{01}\\
y_{01}
\end{pmatrix}=0"

"x_{01}=1"

"y_{01}=2"


for "\lambda_2" :

"\begin{pmatrix}
-2 & -1\\
-4 & -2
\end{pmatrix}
\begin{pmatrix}
x_{02}\\
y_{02}
\end{pmatrix}=0"

"x_{02}=1"

"y_{02}=-2"


Common solution:

"\begin{pmatrix}
x\\
y
\end{pmatrix}=C_1
\begin{pmatrix}
1\\
2
\end{pmatrix}*e^{2t}+C_2
\begin{pmatrix}
1\\
-2
\end{pmatrix}*e^{6t}"


When t=0, x(0)=1 and y(0)=1, so

"\begin{pmatrix}
1\\
1
\end{pmatrix}=C_1
\begin{pmatrix}
1\\
2
\end{pmatrix}+C_2
\begin{pmatrix}
1\\
-2
\end{pmatrix}"


Solution: "C_1=\frac{3}{4}" "C_2=\frac{1}{4}"


Solution of the system:

"\begin{pmatrix}
x\\
y
\end{pmatrix}=\frac{3}{4}
\begin{pmatrix}
1\\
2
\end{pmatrix}*e^{2t}+\frac{1}{4}
\begin{pmatrix}
1\\
-2
\end{pmatrix}*e^{6t}"